# Ground-Compose-Reinforce: Tasking Reinforcement Learning Agents through Formal Language

### v1

@@ A.2 Approximating $V^{*}_{\mathcal{R}}(s,u)$ with Rewarding Self-Loop Transitions
For the previous lava example, the expected return for the option that corresponds to the transition labelled $\langle\lnot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}lava},0\rangle$ now correctly reflects a reward of $-1$ obtained for each step until the option terminates by the agent exiting the lava. We modify our approximation of ${V}^{*}_{\mathcal{R}}(s,u)$ as follows:
@@
6:     end for
7:  end while
8:  return  $v^{*}_{\mathcal{R}}$
A.3 Sources of Approximation Errors
@@
A.4 Training Primitive Value Functions
PVFs can be trained directly from the grounding dataset $\mathcal{D}$ based on any offline RL approach. Algorithm 4 shows how to train the PVF for a single primitive task $\Diamond x$ using a simple offline RL algorithm (Fitted Q-Iteration [70]). The approach can be easily adapted to negations $\Diamond\lnot x$ as well, and in practice, we simultaneously train all $2|\mathcal{AP}|$ possible PVFs in parallel as a single neural network.
@@ A.4 Training Primitive Value Functions
9:  end while
10:  return  $V_{\psi}$
A.5 Potential-based Reward Shaping
@@
0:  MDP $\mathcal{M}$ without rewards, Propositional symbols $\mathcal{AP}$, Dataset $\mathcal{D}$ of labelled trajectories, RM task $\mathcal{R}$ over $\mathcal{AP}$, Shaping potential weighting coefficient $\lambda$ // Phase 1: Ground symbols via offline learning
1:  Train labelling function ^$\hat{\mathcal{L}}(s)$ on $\mathcal{D}$ using any binary classifier
2:  Train PVFs $V^{*}_{\Diamond x}(s)$ and $V^{*}_{\Diamond\lnot x}(s)$, $\forall x\in\mathcal{AP}$ on $\mathcal{D}$ // Phase 2: Solve any RM task using self-generated rewards
@@
11:        Update RM: $u^{\prime}\leftarrow\delta_{u}(u,\hat{\omega})$, $r\leftarrow\delta_{r}(u,\hat{\omega})$
@@ B.1 Domain Descriptions
We considered two novel environments. GeoGrid is an $8\times 8$ image-based gridworld depicted in Figure 1 with six objects randomly positioned at the start of each episode. States are $8\times 8\times 6$-dimensional images that identify each cell’s colour/shape and the agent’s location, while propositions $\{\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}},\mathrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}G}},\mathrm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}},\triangle,\Circle\}$ identify if the agent is at an object with that particular colour or shape. DrawerWorld is a MuJoCo environment adapted from MetaWorld [85]. The agent controls a robotic gripper and can interact with two drawers (left and right) and three boxes (red, green, and blue). Observations are 78-dimensional vectors representing positions of objects and the gripper. Propositions identify whether a particular drawer is open, whether a particular block is picked up by the agent, and whether a particular block is currently inside a particular drawer.
@@
We carefully curated datasets $\mathcal{D}$ in each environment to support our analysis of compositional generalization (RQ2). In GeoGrid, $\mathcal{D}$ is comprised of 5000 trajectories of length 100 generated under a random-action policy. In DrawerWorld, $\mathcal{D}$ is comprised of 350 trajectories of varying length that we collected by manually controlling the robot in the simulator. To ensure sufficient state coverage in $\mathcal{D}$, drawers and boxes were initialized in a random configuration when collecting each trajectory. Behaviours that appear in the dataset include opening and closing drawers, picking up boxes, and moving boxes from one location to another, but no trajectory involves direct interaction with more than one box. A small number of trajectories involve incidental (but not prolonged) interaction with more than one box (e.g. bumping into one box while moving another).
@@
We designed a diverse set of RM tasks (Table 1). For tasks that involve achieving a (temporally extended) goal, the RM terminates and provides a reward of 1 upon doing so. The RM terminates with a reward of 0 if the goal becomes logically impossible (e.g. due to breaking a constraint), except in Safety, which terminates with a penalty of $-1$. Loop and Lift-Red-Box involve repeating some desired behaviour, and the RM yields a reward of 1 for each such repetition.
@@ B.2 Baseline Implementation Details
GeoGrid (propositions are named r,g,b,c,t instead of $\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}},\mathrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}G}},\mathrm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}},\triangle,\Circle$ to avoid confusion with LTL operators)
@@
Lift-Red-Box $\Box\Diamond\mathrm{RedBoxLifted}$
@@
[…]
@@ B.3 Experimental Setup and Hyperparameter Details
All experiments were run on a compute cluster. Supervised training on $\mathcal{D}$ required a single GPU and CPU, minimal memory resources (24GB of RAM or less) and no more than 30 minutes to train any method to 100 epochs. We tuned hyperparameters via a line search over batch size, learning rate, L1 regularization coefficient, and epochs (in that order) using a held-out 10% of the trajectories in D, and the final hyperparameters are reported in Table 5. Final models were retrained on the full data with the tuned hyperparameteres. We note that the batch size hyperparameter should be interpreted differently for methods requiring a GRU. For GCR, it refers to the number of transitions sampled from $\mathcal{D}$. For LTL-BC, Bespoke Reward Model and Bespoke Behaviour Cloning, it refers to the number of full length trajectories sampled from $\mathcal{D}$. This is because it is necessary to keep transitions in a trajectory in the correct order to train the GRU.
Details: RL Training.
B.4 Learning Curves
We report RL learning curves for each method and task in Figures 5 and 6. GCR and its variants have an internal reward model that is better aligned with the ground truth compared to Bespoke Reward Model. In GeoGrid, GCR’s reward model is near perfect, and in all cases, optimizing its internal rewards improves ground truth performance as well. Bespoke Reward Model almost always predicts near-zero rewards on GeoGrid tasks (except Safety), likely since there are few examples of positive demonstrations in $\mathcal{D}$. On DrawerWorld, Bespoke Reward-Model is highly misaligned, predicting large rewards but garnering near-zero return based on the ground truth.
In terms of sample efficiency, we observe that GCR outperforms all baselines. The difference is marginal in GeoGrid, where exploration is less of an issue, but in DrawerWorld, all other reward shaping approaches fail.
Appendix C Autoformalizing Natural Language to Reward Machines
While several works have been dedicated to the autoformalization of natural language instructions into formal languages such as LTL [34, 35, 36, 37], we show that a modern large language model can perform this task zero-shot for RMs, without having specifically been trained on it (to the best of our knowledge). This allows us to directly task the RL agent in our framework through natural language, autoformalize the task description into an RM, and train the agent using its grounded formal language understanding. Critically, the autoformalizer’s understanding of language does not need to be grounded.
ChatGPT-4o correctly produced RMs for all tasks, except for Logic, which requires a complex RM (our solution involved 10 states and 18 transitions). However, we note that ChatGPT-4o was nearly correct for all five trials for Logic—each of its outputted RMs deviated by a single transition that changed the behaviour of the resultant reward function. o3 outputted correct RMs on all tasks. Notably, the outputted RMs were identical in structure to the intended RMs we manually constructed in all cases (with the only differences being in the numbering of RM states and the representation of equivalent logical formulas).
[…]
Environment Description: The environment is a gridworld, where some squares have objects. Each object has a colour (red, green, or blue) and a shape (circle, or triangle).
Propositions:
- red: The agents current cell has a red object.
- blue: The agents current cell has a blue object.
- green: The agents current cell has a green object.
- triangle: The agents current cell has a triangle.
- circle: The agents current cell has a circle.
Task: <TASK DESCRIPTION>
Listing 1: Reward Machine Autoformalization Prompt
Task Description GPT-4o Success Rate o3 Success Rate
Appendix D Societal Impact
Data-efficient learning lowers environmental cost. Ground-Compose-Reinforce (GCR) achieves strong generalization from a relatively small, task-agnostic trajectory dataset. Because it avoids the need for internet-scale demonstrations, the total compute and data collection burden is substantially reduced, which in turn diminishes the carbon footprint of training and retraining large decision-making systems.
Transparent, verifiable task specifications. By exposing an explicit formal specification layer (Reward Machines) between the human and the agent, GCR allows auditors to read, simulate, and formally verify the reward logic before deployment. This contrasts with opaque end-to-end reward models and can help regulators trace undesirable behaviour back to a concrete symbolic condition rather than a latent neural representation, supporting safer and more accountable RL pipelines.
Broader access to capable agents. Because symbols are grounded once and then recomposed to create an unbounded task space, domain experts without ML backgrounds can author complex tasks simply by writing formulas, potentially democratizing advanced robotics and simulation tools in education, manufacturing, and assistive settings. The same mechanism lets small-lab researchers prototype complex multi-stage tasks without the costs associated with collecting new labelled rewards.
Reward hacking and specification gaps. If the learned grounding model assigns the wrong propositions to some states, the formal specification no longer matches human intent. In the paper, we caution that such mis-grounding can lead to “unintended or harmful behaviours” despite the use of a precise formal language. Attackers could also deliberately craft adversarial inputs that flip propositional truths so that the agent believes it has satisfied (or avoided) certain sub-goals, creating safety hazards.
Misuse through automated physical systems. Because GCR enables non-experts to compose lengthy, conditional task logic, it could accelerate deployment of autonomous robots in manufacturing, warehousing, or surveillance. In the absence of stringent safety checks, these robots might execute hazardous motion plans or be repurposed for malign objectives (e.g., automating the handling of dangerous substances or aiding illicit surveillance).
Labour displacement. Easier programming of general-purpose robotic skills may substitute for manual labour in logistics or assembly lines, contributing to job displacement without adequate social safety nets.
@@ Appendix D Societal Impact
Shaping potential weighting coefficient ($\lambda$) 0.1 in Lift-Red-Box, otherwise 1
GCR LTL-BC Bespoke Reward Model Bespoke BC
Labelling Function PVFs
GeoGrid
DrawerWorld
Table 8: Network Architectures for PPO.
GCR Bespoke Reward Model
GeoGrid
Encoder: Conv2d(6,16,3,1,1) ReLU Conv2d(16,32,3,1,1) ReLU Flatten Actor Head: Linear(2048 + $|\mathcal{U}|$,128) ReLU Linear(128,64) ReLU Linear(64,4) Critic Head: Linear(2048 + $|\mathcal{U}|$,128) ReLU Linear(128,64) ReLU Linear(64,1) Encoder: Conv2d(6,16,3,1,1) ReLU Conv2d(16,32,3,1,1) ReLU Flatten GRU(2048,128) Actor Head: Linear(128,128) ReLU Linear(128,64) ReLU Linear(64,4) Critic Head: Linear(128,128) ReLU Linear(128,64) ReLU Linear(64,1)
DrawerWorld
Actor: Linear(78 + $|\mathcal{U}|$,512) ReLU Linear(512,512) ReLU Linear(512,512) ReLU Linear(512,8) Critic: Linear(78 + $|\mathcal{U}|$,512) ReLU Linear(512,512) ReLU Linear(512,512) ReLU Linear(512,1) Encoder: Linear(78,512) ReLU Linear(512,512) GRU(512,512) Actor Head: Linear(512,512) ReLU Linear(512,8) Critic Head: Linear(512,512) ReLU Linear(512,1)
Figure 5: RL learning curves for GeoGrid tasks, showing returns under the agent’s own reward model (“perceived rewards”) as well as rewards under the ground-truth $\mathcal{L}^{*}$ (“actual rewards”). Error bars show standard error.
Figure 6: RL learning curves for DrawerWorld tasks, showing returns under the agent’s own reward model (“perceived rewards”) as well as rewards under the ground-truth $\mathcal{L}^{*}$ (“actual rewards”). Error bars show standard error.

### v2

@@ A.3 Approximating $V^{*}_{\mathcal{R}}(s,u)$ with Rewarding Self-Loop Transitions
For the previous lava example, the expected return for the option that corresponds to the transition labelled $\langle\lnot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}lava},0\rangle$ now correctly reflects a reward of 1-1 obtained for each step until the option terminates by the agent exiting the lava. We modify our approximation of ${V}^{*}_{\mathcal{R}}(s,u)$ as follows:
@@
6:return $v^{*}_{\mathcal{R}}$
A.4 Sources of Approximation Errors
@@
A.5 Training Primitive Value Functions
PVFs can be trained directly from the trajectory dataset $\mathcal{D}$ based on any offline RL approach. Algorithm 4 shows how to train the PVF for a single primitive task $\Diamond x$ using a simple offline RL algorithm (Fitted Q-Iteration [85]). The approach can be easily adapted to negations $\Diamond\lnot x$ as well, and in practice, we simultaneously train all $2|\mathcal{AP}|$ possible PVFs in parallel as a single neural network.
@@ A.5 Training Primitive Value Functions
9:return $V_{\psi}$
A.6 Potential-based Reward Shaping
@@
0: MDP $\mathcal{M}$ without rewards, Propositional symbols $\mathcal{AP}$, Dataset $\mathcal{D}$ of labelled trajectories, RM task $\mathcal{R}$ over $\mathcal{AP}$, Shaping potential weighting coefficient $\lambda$ // Pretraining phase
1: Train labelling function $\hat{\mathcal{L}}(s)$ on $\mathcal{D}$ using any binary classification method
2:Train PVFs $V^{*}_{\Diamond x}(s)$ and $V^{*}_{\Diamond\lnot x}(s)$, $\forall x\in\mathcal{AP}$ on $\mathcal{D}$ // Behaviour elicitation phase
@@
[…]
@@ B.1 Domain Descriptions
[…]
@@
We carefully curated datasets $\mathcal{D}$ in each environment to support our analysis of compositional generalization (RQ2). In GeoGrid, $\mathcal{D}$ is comprised of 5000 trajectories of length 100 generated under a random-action policy. In DrawerWorld, $\mathcal{D}$ is comprised of 350 trajectories of varying length that we collected by manually controlling the robot in the simulator. To ensure sufficient state coverage in $\mathcal{D}$, drawers and boxes were initialized in a random configuration when collecting each trajectory. Behaviours that appear in the dataset include opening and closing drawers, picking up boxes, and moving boxes from one location to another, but no trajectory involves direct interaction with more than one box. A small number of trajectories involve incidental (but not prolonged) interaction with more than one box (e.g. bumping into one box while moving another). We intentionally include accidental behaviours in the DrawerWorld dataset such as failing to grip a box, dropping a box while attempting to move it, and opening a drawer beyond its limit. We also include behaviours not tied to downstream tasks such as placing a box on top of a drawer or throwing a box off the table.
@@
We designed a diverse set of RM tasks (Table 1). For tasks that involve achieving a (temporally extended) goal, the RM terminates and provides a reward of 1 upon doing so. The RM terminates with a reward of 0 if the goal becomes logically impossible (e.g. due to breaking a constraint), except in Safety, which terminates with a penalty of 1-1. Loop and Hold-Red-Box involve repeating some desired behaviour, and the RM yields a reward of 1 for each such repetition. For the precise encodings of tasks as RMs, please see the released code.
@@ B.2 Baseline Implementation Details
GCR LTL-BC Bespoke Reward Model Bespoke BC
Labelling Function PVFs
GeoGrid
DrawerWorld
@@
In GeoGrid, PVFs were trained using Fitted Q-Iteration [85]. In DrawerWorld, PVFs were trained to directly predict Monte Carlo returns. We also considered a state-of-the-art offline RL method, MCQ [86], but it performed worse than Monte Carlo regression. We attribute this to the relatively small size of $\mathcal{D}$ compared to standard offline RL benchmarks.
@@
Table 4: Network Architectures for PPO.
GeoGrid DrawerWorld
GCR Bespoke Reward Model GCR Bespoke Reward Model
[…]
@@
[…]
@@
Hold-Red-Box $\Box\Diamond\mathrm{RedBoxLifted}$
@@
Show-Green-Box $\neg[\neg(\mathrm{GreenBoxInDrawer1}\wedge\Diamond(\mathrm{Drawer1Open}\,\mathcal{U}\,\mathrm{GreenBoxLifted}))$
   $\wedge\neg(\mathrm{GreenBoxInDrawer2}\wedge\Diamond(\mathrm{Drawer2Open}\,\mathcal{U}\,\mathrm{GreenBoxLifted}))$
   $\wedge\neg(\neg\mathrm{GreenBoxInDrawer1}\wedge\neg\mathrm{GreenBoxinDrawer2}\wedge\Diamond\mathrm{GreenBoxLifted})]$
@@ B.3 Experimental Setup and Hyperparameter Details
All experiments were run on a compute cluster. Supervised training on $\mathcal{D}$ required a single GPU and CPU, minimal memory resources (24GB of RAM or less) and no more than 30 minutes to train any method to 100 epochs. We tuned hyperparameters via a line search over batch size, learning rate, L1 regularization coefficient, and epochs (in that order) using a held-out 10% of the trajectories in D, and the final hyperparameters are reported in Table 6. Final models were retrained on the full data with the tuned hyperparameters. We note that the batch size hyperparameter should be interpreted differently for methods requiring a GRU. For GCR, it refers to the number of transitions sampled from $\mathcal{D}$. For LTL-BC, Bespoke Reward Model and Bespoke Behaviour Cloning, it refers to the number of full length trajectories sampled from $\mathcal{D}$. This is because it is necessary to keep transitions in a trajectory in the correct order to train the GRU.
@@
Shaping potential weighting coefficient ($\lambda$) 0.1 in Hold-Red-Box, otherwise 1
B.4 Learning Curves
Figure 6: RL learning curves for GeoGrid, showing returns under the agent’s own reward model (“perceived rewards”) and under the ground-truth $\mathcal{L}^{*}$ (“actual rewards”). Perceived rewards are reported without shaped rewards, and shaded regions show standard error. Approaches based on Ground-Compose-Reinforce (including No RS and High-Level RS) generate rewards that are closely aligned with the ground truth and lead to an effective final policy, while Bespoke Reward Model produces near-zero rewards in all cases and makes little progress.
Figure 7: RL learning curves for DrawerWorld, showing returns under the agent’s own reward model (“perceived rewards”) and under the ground-truth $\mathcal{L}^{*}$ (“actual rewards”). Perceived rewards are reported without shaped rewards, and shaded regions show standard error. When evaluated under ground-truth rewards, Ground-Compose-Reinforce with our reward shaping strategy learns strong policies in all cases, while alternative approaches make no progress. Notably, Bespoke Reward Model results in a final policy with high perceived rewards, but poor actual performance.
We report RL learning curves for each method and task in Figures 6 and 7. GCR and its variants have an internal reward model that is better aligned with the ground truth compared to Bespoke Reward Model. In GeoGrid, GCR’s reward model is near perfect, and in all cases, optimizing its internal rewards improves ground truth performance as well. Bespoke Reward Model almost always predicts near-zero rewards on GeoGrid tasks (except Safety), likely since there are few examples of positive demonstrations in $\mathcal{D}$. On DrawerWorld, Bespoke Reward-Model is highly misaligned, predicting large rewards but garnering near-zero return based on the ground truth.
In terms of sample efficiency, we observe that GCR with our reward shaping strategy outperforms all baselines. The difference is marginal in GeoGrid, where exploration is less of an issue, but in DrawerWorld, all other reward shaping approaches fail.
Appendix C Autoformalizing Natural Language to Reward Machines
While several works have been dedicated to the autoformalization of natural language instructions into formal languages such as LTL [36, 37, 38, 39], we show that a modern LLM can sometimes perform this task zero-shot for RMs, without having specifically been trained on it (to the best of our knowledge). This allows us to directly task the RL agent in our framework through natural language, then autoformalize the task description into an RM.
ChatGPT-4o correctly produced RMs for all tasks, except for Logic, which requires a complex RM (our solution involved 10 states and 18 transitions). However, we note that ChatGPT-4o was nearly correct for all five trials for Logic—each of its outputted RMs deviated by a single transition that changed the behaviour of the resultant reward function. o3 outputted correct RMs on all tasks. Notably, the outputted RMs were identical in structure to the intended RMs we manually constructed in all cases (with the only differences being in the naming of RM states and the representation of equivalent logical formulas).
[…]
Environment Description: The environment is a gridworld, where some squares have objects. Each object has a single colour (red, green, or blue) and a single shape (circle, or triangle).
Propositions:
- red: The agents current cell has a red object.
- blue: The agents current cell has a blue object.
- green: The agents current cell has a green object.
- triangle: The agents current cell has a triangle.
- circle: The agents current cell has a circle.
Task: <TASK DESCRIPTION>
Listing 1: Reward Machine Autoformalization Prompt
Task Description GPT-4o Success Rate o3 Success Rate
Appendix D Societal Impact
Data‑efficient learning lowers environmental cost. Ground‑Compose‑Reinforce (GCR) achieves strong generalization from a relatively small, task‑agnostic trajectory dataset. Because it avoids the need for internet‑scale demonstrations, the total compute and data collection burden is substantially reduced, which in turn diminishes the carbon footprint of training and retraining large decision‑making systems.
Transparent, verifiable task specifications. By exposing an explicit formal specification layer (Reward Machines) between the human and the agent, GCR allows auditors to read, simulate, and formally verify the reward logic before deployment. This contrasts with opaque end‑to‑end reward models and can help regulators trace undesirable behaviour back to a concrete symbolic condition rather than a latent neural representation, supporting safer and more accountable RL pipelines.
Broader access to capable agents. Because symbols are grounded once and then recomposed to create an unbounded task space, domain experts without ML backgrounds can author complex tasks simply by writing RMs, potentially democratizing advanced robotics and simulation tools in education, manufacturing, and assistive settings. The same mechanism lets small‑lab researchers prototype complex multi‑stage tasks without the costs associated with collecting new labelled rewards.
Reward hacking and specification gaps. If the learned interpretation of propositions in the environment is erroneous, the resultant behaviour may no longer match human intent. In the paper, we caution that such mis-grounding can lead to harmful behaviours despite the use of a precise formal specification.
Labour displacement. Easier programming of general‑purpose robotic skills may substitute for manual labour in logistics or assembly lines, contributing to job displacement without adequate social safety nets.